OpenStudy (anonymous):

if i drive 50 mph and speed up steadily to 70mph over a 30 min period, how far will i have traveled during the 30 min period of acceleration?

OpenStudy (lastdaywork):

First find acceleration then use - s = ut + (1/2)at^2 The above ^^ analysis assumes that you don't crash :P

OpenStudy (anonymous):

i have been ttying but can't figure this problem out. i am getting so frustrated, i have an exam tomorrow and don't want to fail

OpenStudy (lastdaywork):

Can you find acceleration ??

OpenStudy (anonymous):

not really i am not good with math

OpenStudy (lastdaywork):

a = (v-u)/t where v = final velocity u = initial velocity t = time period

OpenStudy (anonymous):

i've been doing speed=x time. i have been racking my brains. can't come up with an ans. according to my practice book, the ans is either 20 miles, 27.5 miles, 35 miles or 42.5 miles. i can't get any of these ans

OpenStudy (lastdaywork):

Find acceleration first; then use the equation - distance = u*t + (1/2)*a*t*t Do one step at a time..

OpenStudy (anonymous):

ok. i will try again but doubt it. thanks

OpenStudy (anonymous):

i am stumped, but thanks anyway have a good niite